Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year: Molding Fabrication Total Machine-hours 20,000 30,000 50,000 Fixed manufacturing overhead costs $ 700,000 $ 210,000 $ 910,000 Variable manufacturing overhead cost per machine-hour $ 3.00 $ 1.00

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs: Job D-70: Molding Fabrication Total Direct materials cost $ 375,000 $ 325,000 $ 700,000 Direct labor cost $ 200,000 $ 160,000 $ 360,000 Machine-hours 14,000 6,000 20,000

Job C-200: Molding Fabrication Total Direct materials cost $ 300,000 $ 250,000 $ 550,000 Direct labor cost $ 175,000 $ 225,000 $ 400,000 Machine-hours 6,000 24,000 30,000

2. Assume Delph uses departmental predetermined overhead rates based on machine hours.

a. Compute the departmental predetermined overhead rates

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200

c. If Delph establishes big prices that are 150% of total manufacturing costs, what bid prices would it have to establish?

d. What is Delph's COGS?
Answer #1

A. Computation of departmental pre determined OH rates

Particulars molding Fabrication
Fixed manufacturing overhead cost per hour 700000/20000=$35 210000/30000=$7
Variable manufacturing overhead cost per hour $3 $1

Total

$38 $8

B. Manufacturing cost of job D 70 and job C-200

Job D-70. Job C-200

Direct material cost. 700000. 550000

Direct labour cost. 360000. 400000

Overhead cost. 760000. 240000

(20000*38). (30000*8)

B.Total.   1830000.   1190000

C. Bid prices shall be(150% of B) 2745000. 1785000

D. COGS = 1830000+1190000. .

=3020000.